OpenStudy (anonymous):

Need to find the indefinite integral of:: Sqrt(1-x^2)dx ** Any way to do that without trig substitution? ** [Tried integrating by parts, did not work]

OpenStudy (experimentx):

try trig sub

OpenStudy (anonymous):

spure it off

OpenStudy (anonymous):

without trig sub.

OpenStudy (experimentx):

are you expecting it to solve without trig sub?

OpenStudy (anonymous):

The thing is, that our lectorer gave us this question before we learnt trig subs. (Maybe by mistake?).

OpenStudy (experimentx):

there might be other ways .. but I don't know yet.

OpenStudy (anonymous):

I=∫▒〖√((1-x^2 ) ) dx〗=∫▒√((1-x^2 ) )×1dx Integrate by parts =√((1-x^2 ) ) x-∫▒d/dx √((1-x^2 ) ) xdx =x√(1-x^2 )-∫▒〖(-2x)/(2√(1-x^2 )) xdx〗 =x√(1-x^2 ) -∫▒(1-x^2-1)/√(1-x^2 ) dx =x√(1-x^2 ) -∫▒√(1-x^2 ) dx+∫▒〖1/√(1-x^2 ) dx〗 =x√(1-x^2 ) –I +sin^(-1)⁡x+c 2I=x√(1-x^2 )+ sin^(-1)⁡x+c I=(x√(1-x^2 ))/2+1/2 sin^(-1)⁡x+C

OpenStudy (experimentx):

interesting ... this is just opposite of differentiation.